Jefferson Lab HALL A Logbook for June 01 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

  • 65201 06/30/01 23:48 solvignon PR left #18
  • 65200 06/30/01 23:41 slifer Shift summary
  • 65199 06/30/01 23:05 Solvignon beam helicity
  • 65196 06/30/01 22:41 apar Start_of_Run_Parity_1057
  • 65193 06/30/01 22:37 slifer beam-time accounting program
  • 65192 06/30/01 21:59 slifer helicity
  • 65191 06/30/01 21:39 W. Korsch Another loose cable!!
  • 65188 06/30/01 21:29 apar Start_of_Run_Parity_1056
  • 65184 06/30/01 21:28 apar Start_of_Run_Parity_1055
  • 65183 06/30/01 21:26 K. McCormick Synchronization of 3 DAQs back?
  • 65178 06/30/01 21:04 apar Start_of_Run_Parity_1054
  • 65173 06/30/01 21:02 apar Start_of_Run_Parity_1053
  • 65168 06/30/01 20:28 apar Start_of_Run_Parity_1052
  • 65165 06/30/01 20:14 R. Michaels datamon quirk fixed
  • 65162 06/30/01 20:07 apar Start_of_Run_Parity_1051
  • 65159 06/30/01 19:24 Solvignon datamon (right arm)
  • 65156 06/30/01 19:17 apar Start_of_Run_Parity_1050
  • 65153 06/30/01 19:15 Jaideep Sing Laser EPR correction
  • 65152 06/30/01 19:14 slifer datamon (right arm)
  • 65151 06/30/01 18:46 slifer datamon problem (right arm)
  • 65150 06/30/01 18:22 xfzhu, Wolfg EPR polarization:36.7%
  • 65147 06/30/01 18:14 apar Start_of_Run_Parity_1049
  • 65142 06/30/01 17:52 apar Start_of_Run_Parity_1048
  • 65139 06/30/01 17:08 Seonho Choi Stability check of ADC/TDC peak position vs time
  • 65138 06/30/01 17:02 Jaideep Sing Left Arm PR 2nd Layer #17
  • 65135 06/30/01 16:51 apar Start_of_Run_Parity_1047
  • 65132 06/30/01 16:48 Seonho Choi Left Arm PRL1 #12 fixed
  • 65131 06/30/01 16:12 Seonho Choi Shift Summary 6/30/01 Day
  • 65128 06/30/01 15:50 K. McCormick Loose background counters Left arm
  • 65126 06/30/01 15:48 apar Start_of_Run_Parity_1046
  • 65122 06/30/01 15:44 Seonho Choi Synchronization of 3 DAQ's
  • 65121 06/30/01 15:32 Jiang,chudak pion rejector layer-1 block-12 fixed
  • 65110 06/30/01 13:09 apar Start_of_Run_Parity_1045
  • 65107 06/30/01 12:58 Wolfgang, Xf Target - EPR - 37 %
  • 65106 06/30/01 12:53 Seonho Choi Hall A BCM and Current Ramping
  • 65103 06/30/01 12:07 apar Start_of_Run_Parity_1044
  • 65100 06/30/01 11:57 Seonho Choi Left Arm PRL1 problem
  • 65099 06/30/01 11:50 Nikolai Savv PR layer 1 #12 ADC problem
  • 65096 06/30/01 10:57 apar Start_of_Run_Parity_1043
  • 65093 06/30/01 10:40 Seonho Choi Hall A BCM & Beam Current Ramping
  • 65090 06/30/01 09:50 apar Start_of_Run_Parity_1042
  • 65085 06/30/01 09:01 apar Start_of_Run_Parity_1041
  • 65082 06/30/01 08:01 kramerk Shift Summary
  • 65080 06/30/01 07:44 kramerk Beam back!
  • 65078 06/30/01 07:43 apar Start_of_Run_Parity_1040
  • 65076 06/30/01 06:00 adaq Runsummary_Left_HRS
  • 65075 06/30/01 06:00 adaq Runsummary_Right_HRS
  • 65074 06/30/01 06:00 X. Zheng Target - EPR - 35.5%
  • 65073 06/30/01 05:52 kramerk Beam down for another hour
  • 65072 06/30/01 05:19 kramerk Target - NMR Polarization - 37.5%
  • 65068 06/30/01 04:23 apar Start_of_Run_Parity_1039
  • 65063 06/30/01 03:21 apar Start_of_Run_Parity_1038
  • 65059 06/30/01 02:49 kramerk Beam back at 10 uAmps
  • 65058 06/30/01 02:46 kramerk Beam Current Low
  • 65056 06/30/01 02:20 apar Start_of_Run_Parity_1037
  • 65051 06/30/01 01:19 apar Start_of_Run_Parity_1036
  • 65047 06/30/01 00:43 Kevin, Wolfg Target - EPR - 38.0 %
  • 65045 06/30/01 00:18 apar Start_of_Run_Parity_1035
  • 65041 06/30/01 00:03 Sulkosky Target-EPR Measurement
  • 65039 06/30/01 00:00 J. Roche Shift Summary
  • Friday

  • 65035 06/29/01 23:16 apar Start_of_Run_Parity_1034
  • 65030 06/29/01 22:14 apar Start_of_Run_Parity_1033
  • 65025 06/29/01 21:16 apar Start_of_Run_Parity_1032
  • 65023 06/29/01 21:09 jroche charge as. : what to do if it's no working
  • 65020 06/29/01 20:15 apar Start_of_Run_Parity_1031
  • 65019 06/29/01 19:33 Sulkosky Target-EPR and Movement
  • 65018 06/29/01 19:31 X. Zheng, Wo Target - EPR - 37.5 %
  • 65012 06/29/01 19:13 apar Start_of_Run_Parity_1030
  • 65008 06/29/01 18:32 apar Start_of_Run_Parity_1029
  • 65003 06/29/01 18:14 apar Start_of_Run_Parity_1028
  • 64998 06/29/01 17:59 apar Start_of_Run_Parity_1027
  • 64993 06/29/01 17:31 apar Start_of_Run_Parity_1026
  • 64990 06/29/01 16:03 Saha Shift Summary
  • 64988 06/29/01 14:36 averett beam RF problems
  • 64987 06/29/01 14:33 Armstrong feedback off
  • 64986 06/29/01 14:01 averett no beam-RF problems
  • 64983 06/29/01 13:58 apar Start_of_Run_Parity_1025
  • 64982 06/29/01 12:01 Antonin Test with PARITY DAQ
  • 64981 06/29/01 11:53 apar Start_of_Run_Parity_1024
  • 64980 06/29/01 11:36 apar Start_of_Run_Parity_1023
  • 64979 06/29/01 10:59 bogdanw NMR result
  • 64978 06/29/01 09:29 X. Zheng Target - EPR - 41.5%
  • 64977 06/29/01 09:13 K. McCormick Number of Events o,+
  • 64975 06/29/01 08:00 M. Rvachev Shift summary
  • 64969 06/29/01 07:14 apar Start_of_Run_Parity_1022
  • 64964 06/29/01 06:10 apar Start_of_Run_Parity_1021
  • 64962 06/29/01 06:00 adaq Runsummary_Left_HRS
  • 64961 06/29/01 06:00 adaq Runsummary_Right_HRS
  • 64960 06/29/01 05:32 xfzhu NMR measurement
  • 64959 06/29/01 05:11 X. Zheng Target - EPR - 41.5%
  • 64957 06/29/01 05:10 apar Start_of_Run_Parity_1020
  • 64945 06/29/01 02:31 apar Start_of_Run_Parity_1018
  • 64941 06/29/01 01:31 apar Start_of_Run_Parity_1017
  • 64936 06/29/01 00:29 apar Start_of_Run_Parity_1016
  • 64932 06/29/01 00:00 Allyn Swing Shift Summary
  • Thursday

  • 64931 06/28/01 23:58 Sulkosky Target-Polarization Measurements: 42 %
  • 64930 06/28/01 23:34 X. Zheng Correction for EPR result
  • 64928 06/28/01 23:28 apar Start_of_Run_Parity_1015
  • 64925 06/28/01 23:26 X. Zheng Target - EPR - 41.94%
  • 64922 06/28/01 22:18 Allyn IOC probelms: current not read out
  • 64919 06/28/01 21:50 apar Start_of_Run_Parity_1014
  • 64914 06/28/01 20:27 Sulkosky Target-Rotate to 270 degrees
  • 64911 06/28/01 20:23 apar Start_of_Run_Parity_1013
  • 64908 06/28/01 20:05 apar Start_of_Run_Parity_1012
  • 64907 06/28/01 20:05 Higinbotham Arc Energy BDL Measurement
  • 64906 06/28/01 19:31 Sarah Magnets
  • 64905 06/28/01 19:19 Sulkosky Target-Move to Empty target for ARC Measurement.
  • 64904 06/28/01 18:17 Sarah Magnet Recovery
  • 64903 06/28/01 18:17 Sulkosky Target-NMR Polarization.
  • 64902 06/28/01 18:03 Wolfgang Target - EPR
  • 64901 06/28/01 18:01 Sulkosky Target-Polarization Measurements
  • 64900 06/28/01 17:52 x. jiang Fixed the software bug in the strange tdc spectra
  • 64899 06/28/01 15:20 Rikki shift summary
  • 64894 06/28/01 13:31 apar Start_of_Run_Parity_1011
  • 64892 06/28/01 11:24 Wolfgang, Xf Target - EPR - 43 %
  • 64888 06/28/01 11:07 apar Start_of_Run_Parity_1010
  • 64886 06/28/01 10:09 Seonho Choi Beam Energy Stability
  • 64882 06/28/01 09:38 kramerk Target - NMR Polarization @ 3:55am - 43.1%
  • 64881 06/28/01 09:22 apar Start_of_Run_Parity_1009
  • 64875 06/28/01 08:00 M. Rvachev Shift summary
  • 64873 06/28/01 07:56 apar Start_of_Run_Parity_1008
  • 64868 06/28/01 06:50 apar Start_of_Run_Parity_1007
  • 64864 06/28/01 06:00 adaq Runsummary_Left_HRS
  • 64863 06/28/01 06:00 adaq Runsummary_Right_HRS
  • 64862 06/28/01 05:49 apar Start_of_Run_Parity_1006
  • 64859 06/28/01 05:19 xfzhu EPR/NMR measurement
  • 64856 06/28/01 04:43 apar Start_of_Run_Parity_1005
  • 64854 06/28/01 03:54 M. Rvachev Left arm CODA reboot
  • 64847 06/28/01 02:45 apar Start_of_Run_Parity_1004
  • 64843 06/28/01 01:47 apar Start_of_Run_Parity_1003
  • 64839 06/28/01 00:45 apar Start_of_Run_Parity_1002
  • 64836 06/28/01 00:25 xfzhu EPR fitting result
  • 64833 06/28/01 00:02 Reitz shift summary
  • 64832 06/28/01 00:00 jpchen Please record target holding field values
  • Wednesday

  • 64831 06/27/01 23:58 Reitz automatic feedback disabled
  • 64829 06/27/01 23:42 apar Start_of_Run_Parity_1001
  • 64826 06/27/01 23:12 jpchen NMR 42% (fit), EPR ~ 41% (eye)
  • 64824 06/27/01 22:53 apar Start_of_Run_Parity_1000
  • 64821 06/27/01 21:44 kramerk Target - Today's NMRs
  • 64820 06/27/01 21:42 R. Michaels re: ped suppression woes
  • 64819 06/27/01 21:29 jpchen beam back
  • 64816 06/27/01 21:02 apar Start_of_Run_Parity_999
  • 64813 06/27/01 20:36 jpchen Beam unstable, and now Beam off to tune for Hall B
  • 64809 06/27/01 19:55 apar Start_of_Run_Parity_998
  • 64804 06/27/01 19:31 apar Start_of_Run_Parity_997
  • 64802 06/27/01 19:10 jpchen NMR at 11:20 was about 37%, lost pol due to a failed NMR before this
  • 64799 06/27/01 18:29 apar Start_of_Run_Parity_996
  • 64795 06/27/01 17:35 apar Start_of_Run_Parity_995
  • 64794 06/27/01 17:12 jpchen NMR and EPR, pol~ 39%
  • 64792 06/27/01 16:35 Seonho Choi R-Arm S1,S2 number of hits for each paddle
  • 64791 06/27/01 16:21 adaq shift summary
  • 64786 06/27/01 15:18 apar Start_of_Run_Parity_994
  • 64785 06/27/01 14:04 apar Start_of_Run_Parity_993
  • 64779 06/27/01 12:32 apar Start_of_Run_Parity_992
  • 64775 06/27/01 11:53 Rikki NMR measurement
  • 64774 06/27/01 11:42 apar Start_of_Run_Parity_991
  • 64772 06/27/01 11:37 apar Start_of_Run_Parity_990
  • 64769 06/27/01 10:06 W. Korsch, J Target - EPR - 40 %
  • 64766 06/27/01 09:48 apar Start_of_Run_Parity_989
  • 64765 06/27/01 09:41 kramerk Target - NMR Polarization @ 00:03 27 Jun - 32.7%
  • 64762 06/27/01 08:36 B. Zhang charge asymmetry
  • 64761 06/27/01 08:19 xfzhu right arm scintillators
  • 64758 06/27/01 08:16 apar Start_of_Run_Parity_988
  • 64754 06/27/01 08:00 M. Rvachev Shift summary
  • 64752 06/27/01 07:16 apar Start_of_Run_Parity_987
  • 64751 06/27/01 06:50 B. Moffit Beam downtime
  • 64748 06/27/01 06:00 adaq Runsummary_Left_HRS
  • 64747 06/27/01 06:00 adaq Runsummary_Right_HRS
  • 64746 06/27/01 05:46 apar Start_of_Run_Parity_986
  • 64741 06/27/01 05:04 xfzhu, Xiaoc EPR plorization : 35.8%
  • 64738 06/27/01 04:43 apar Start_of_Run_Parity_985
  • 64737 06/27/01 04:40 Nilanga Left Arm VDC TDC problem: due to software
  • 64736 06/27/01 04:19 M. Rvachev Need for a pedestal run?
  • 64735 06/27/01 04:00 B. Moffit Beam downtime
  • 64732 06/27/01 03:59 xfzhu Problem with right arm scintllators ADC hit pattens
  • 64730 06/27/01 03:06 apar Start_of_Run_Parity_984
  • 64726 06/27/01 02:05 apar Start_of_Run_Parity_983
  • 64724 06/27/01 01:42 B. Moffit Beam downtime
  • 64719 06/27/01 00:52 apar Start_of_Run_Parity_982
  • 64718 06/27/01 00:38 xfzhu , Wolf EPR/NMR , holding field changed to 270 degree
  • 64715 06/27/01 00:15 apar Start_of_Run_Parity_981
  • 64711 06/27/01 00:00 Reitz swing shift summary
  • Tuesday

  • 64710 06/26/01 23:49 Armstrong spot looks good
  • 64707 06/26/01 23:36 apar Start_of_Run_Parity_980
  • 64706 06/26/01 23:32 Armstrong go to BeO target
  • 64705 06/26/01 21:09 Armstrong charge asymmetry over time
  • 64704 06/26/01 20:41 X. Zheng, W. Target - EPR - 32 %
  • 64703 06/26/01 20:32 Saha Arc Energy measurement scrapped
  • 64702 06/26/01 19:55 Saha Arc Integral Measurement
  • 64701 06/26/01 19:18 Armstrong check spot/position with BeO
  • 64700 06/26/01 18:49 Armstrong lasers off, move to BeO
  • 64699 06/26/01 17:50 Armstrong Compton electron detector moved out
  • 64698 06/26/01 17:12 K. McCormick He Target IOC Swapped with spare
  • 64697 06/26/01 16:43 Armstrong sweep begins
  • 64696 06/26/01 16:00 kramerk Target - NMR Measurement - 27.5 %
  • 64692 06/26/01 15:31 jpchen Limit to low current on BeO
  • 64691 06/26/01 15:27 jpchen Online survey numbers
  • 64690 06/26/01 15:24 R. Michaels re: multiple peaks in R-arm S1 pad 1R
  • 64689 06/26/01 15:18 jpchen laser profile check
  • 64688 06/26/01 14:54 atrig Trigger-Download
  • 64676 06/26/01 12:31 jpchen lasers ready to be turned on
  • 64675 06/26/01 12:30 jpchen Air flow and heaters for oven are back on
  • 64674 06/26/01 12:22 jpchen target survey completed
  • 64671 06/26/01 10:57 Antonin Test runs after changing a cable
  • 64670 06/26/01 10:50 apar Start_of_Run_Parity_979
  • 64669 06/26/01 10:48 apar Start_of_Run_Parity_978
  • 64668 06/26/01 10:46 apar Start_of_Run_Parity_977
  • 64667 06/26/01 10:43 apar Start_of_Run_Parity_976
  • 64666 06/26/01 09:43 jpchen Target survey starts, lasers off/locked out, oven off, ari flow off, Cooling jet He flow off
  • 64665 06/26/01 08:53 kramerk Target - NMR Polarization @ 7:47pm Jun 24 - 35.8%
  • 64663 06/26/01 07:37 M. Rvachev Shift summary
  • 64660 06/26/01 06:57 X. Zheng Target - EPR - 36.2%
  • 64658 06/26/01 06:35 apar Start_of_Run_Parity_975
  • 64654 06/26/01 06:01 X. Zheng June 26 summary of electrons @ x=0.6
  • 64653 06/26/01 06:00 adaq Runsummary_Left_HRS
  • 64652 06/26/01 06:00 adaq Runsummary_Right_HRS
  • 64651 06/26/01 05:49 M. Rvachev Beam charge asymmetry oscillations
  • 64650 06/26/01 05:35 apar Start_of_Run_Parity_974
  • 64645 06/26/01 04:50 B.Moffit N2 Cryo Levels
  • 64644 06/26/01 04:36 B. Moffit N2 Cryo Levels
  • 64643 06/26/01 04:31 apar Start_of_Run_Parity_973
  • 64638 06/26/01 03:28 apar Start_of_Run_Parity_972
  • 64633 06/26/01 02:21 apar Start_of_Run_Parity_971
  • 64628 06/26/01 01:05 apar Start_of_Run_Parity_970
  • 64623 06/26/01 00:04 xfzhu Right arm scintillator
  • 64622 06/26/01 00:03 Armstrong/Su ROC 15 disconnect etc.
  • Monday

  • 64620 06/25/01 23:59 K. McCormick 6/25 SWING Summary
  • 64617 06/25/01 23:56 apar Start_of_Run_Parity_969
  • 64611 06/25/01 23:50 apar Start_of_Run_Parity_968
  • 64610 06/25/01 23:43 xfzhu rigth arm scintillator
  • 64608 06/25/01 23:21 Armstrong beam charge asymmetry redux
  • 64607 06/25/01 23:21 xfzhu EPR polarization
  • 64606 06/25/01 23:06 Armstrong Target: holding field decreasing with time
  • 64603 06/25/01 22:55 apar Start_of_Run_Parity_967
  • 64600 06/25/01 21:56 Armstrong ROC 14 disconnect on Left arm
  • 64597 06/25/01 21:51 apar Start_of_Run_Parity_966
  • 64593 06/25/01 21:09 apar Start_of_Run_Parity_965
  • 64590 06/25/01 20:08 Armstrong target - NMR
  • 64587 06/25/01 20:05 apar Start_of_Run_Parity_964
  • 64582 06/25/01 18:32 apar Start_of_Run_Parity_963
  • 64579 06/25/01 17:59 Armstrong Harp Scan in Arc
  • 64578 06/25/01 17:21 apar Start_of_Run_Parity_962
  • 64571 06/25/01 16:18 apar Start_of_Run_Parity_961
  • 64568 06/25/01 16:01 Adaq shift summary
  • 64566 06/25/01 15:59 W. Korsch/jp Transverse gradient field reduced to 5A
  • 64565 06/25/01 15:58 jpchen/W. Ko EPR, pol=35.7%
  • 64562 06/25/01 15:12 apar Start_of_Run_Parity_960
  • 64557 06/25/01 14:13 apar Start_of_Run_Parity_959
  • 64552 06/25/01 13:10 apar Start_of_Run_Parity_958
  • 64547 06/25/01 12:11 apar Start_of_Run_Parity_957
  • 64544 06/25/01 11:12 Binet Solution to the compton problem
  • 64541 06/25/01 11:09 apar Start_of_Run_Parity_956
  • 64538 06/25/01 10:34 R. Michaels hit patterns on R-arm
  • 64537 06/25/01 10:30 jpchen/X. Zh EPR
  • 64534 06/25/01 10:07 apar Start_of_Run_Parity_955
  • 64531 06/25/01 09:03 adaq beam half wave plate
  • 64529 06/25/01 09:00 apar Start_of_Run_Parity_954
  • 64523 06/25/01 08:03 M. Rvachev Shift summary
  • 64522 06/25/01 07:56 X. Zheng summary of electrons for x=0.6, OUT/IN
  • 64521 06/25/01 07:55 M. Rvachev Rebooted hiresxt26
  • 64518 06/25/01 07:44 apar Start_of_Run_Parity_953
  • 64514 06/25/01 06:32 apar Start_of_Run_Parity_952
  • 64510 06/25/01 06:09 M. Rvachev "spec_R_s2_TDC_R_1" continued
  • 64509 06/25/01 06:00 adaq Runsummary_Left_HRS
  • 64508 06/25/01 06:00 adaq Runsummary_Right_HRS
  • 64507 06/25/01 05:32 apar Start_of_Run_Parity_951
  • 64502 06/25/01 05:14 M. Rvachev "spec_R_s2_TDC_R_1" continued
  • 64500 06/25/01 04:34 apar Start_of_Run_Parity_950
  • 64494 06/25/01 03:34 apar Start_of_Run_Parity_949
  • 64491 06/25/01 03:24 moffit Nitrogen espace header
  • 64490 06/25/01 03:09 moffit Nitrogen espace header
  • 64489 06/25/01 02:56 X. Zheng Target - EPR - 35.8%
  • 64487 06/25/01 02:32 apar Start_of_Run_Parity_948
  • 64483 06/25/01 01:35 apar Start_of_Run_Parity_947
  • Sunday

  • 64474 06/24/01 23:59 Hong Xiang Shift Summary
  • 64472 06/24/01 23:41 Armstrong beam oscillations gone?
  • 64470 06/24/01 23:30 apar Start_of_Run_Parity_945
  • 64463 06/24/01 22:50 R. Michaels safety limit on auto chg asy feedback
  • 64460 06/24/01 22:15 apar Start_of_Run_Parity_944
  • 64457 06/24/01 21:45 Armstrong suggestion read asymmetry feedback
  • 64456 06/24/01 21:23 Armstrong reboot adaqcp
  • 64455 06/24/01 21:14 apar Start_of_Run_Parity_943
  • 64446 06/24/01 20:15 Wolfgang Target - EPR - 34.9 %
  • 64443 06/24/01 19:37 apar Start_of_Run_Parity_942
  • 64440 06/24/01 19:09 Armstrong "Hall A Charge Asymmetry" GUI
  • 64439 06/24/01 18:48 Wolfgang Target: Correct Grad. Coil Currents
  • 64438 06/24/01 18:46 Binet Compton polarimetry shift summary
  • 64437 06/24/01 18:29 Armstrong FastFeedback is off (and has been off)
  • 64436 06/24/01 18:27 xfzhu problem with right arm scintililator: spec_R_s2_TDC_R_1
  • 64433 06/24/01 18:24 apar Start_of_Run_Parity_941
  • 64430 06/24/01 18:02 xfzhu Problem with right arm scintililator: spec_R_s2_TDC_R_1
  • 64429 06/24/01 17:54 Armstrong oscillating asymmetry and beam current
  • 64428 06/24/01 17:26 Armstrong Target PC crash
  • 64420 06/24/01 16:13 apar Start_of_Run_Parity_939
  • 64417 06/24/01 16:08 kramerk Target - NMR @ 2:29pm - 36.0%
  • 64415 06/24/01 15:09 apar Start_of_Run_Parity_938
  • 64412 06/24/01 14:41 B. Zhang NMR
  • 64411 06/24/01 14:18 A .Deur Beam down because of cavity problem.
  • 64408 06/24/01 14:06 Wolfgang Change Polarity on Trans. Gradient Coil
  • 64407 06/24/01 14:04 Wolfgang Target - EPR - 37.2 %
  • 64405 06/24/01 13:15 apar Start_of_Run_Parity_937
  • 64400 06/24/01 11:55 apar Start_of_Run_Parity_936
  • 64395 06/24/01 10:54 apar Start_of_Run_Parity_935
  • 64391 06/24/01 10:13 kramerk Target - Revised NMR Polarizations for Jun. 23
  • 64389 06/24/01 09:52 apar Start_of_Run_Parity_934
  • 64384 06/24/01 08:47 apar Start_of_Run_Parity_933
  • 64379 06/24/01 08:00 Hong Xiang Shift Summary
  • 64378 06/24/01 07:43 apar Start_of_Run_Parity_932
  • 64372 06/24/01 07:21 apar Start_of_Run_Parity_931
  • 64370 06/24/01 06:49 X. Zheng Target - EPR - 36.4%
  • 64367 06/24/01 06:03 apar Start_of_Run_Parity_930
  • 64364 06/24/01 06:00 adaq Runsummary_Left_HRS
  • 64363 06/24/01 06:00 adaq Runsummary_Right_HRS
  • 64360 06/24/01 04:56 apar Start_of_Run_Parity_929
  • 64355 06/24/01 03:33 apar Start_of_Run_Parity_928
  • 64350 06/24/01 01:53 apar Start_of_Run_Parity_927
  • 64345 06/24/01 00:44 apar Start_of_Run_Parity_926
  • 64342 06/24/01 00:27 X. Zheng Target - EPR - 36.1%
  • 64338 06/24/01 00:18 apar Start_of_Run_Parity_925
  • 64335 06/24/01 00:14 apar Start_of_Run_Parity_924
  • 64330 06/24/01 00:00 strauch Swing shift summary
  • Saturday

  • 64329 06/23/01 23:39 Binet Compton polarimetry measurement
  • 64325 06/23/01 23:02 apar Start_of_Run_Parity_923
  • 64321 06/23/01 22:01 apar Start_of_Run_Parity_922
  • 64318 06/23/01 21:29 Binet Compton Problem
  • 64315 06/23/01 21:02 apar Start_of_Run_Parity_921
  • 64312 06/23/01 19:12 Binet Compton chicane
  • 64311 06/23/01 18:58 Binet beam in the Compton chicane
  • 64308 06/23/01 18:54 apar Start_of_Run_Parity_920
  • 64306 06/23/01 18:53 E.Chudakov Beam line magnet change
  • 64304 06/23/01 18:37 Binet Compton chicane problem
  • 64303 06/23/01 18:25 xfzhu, Korsc EPR measurement
  • 64300 06/23/01 17:52 apar Start_of_Run_Parity_919
  • 64295 06/23/01 16:46 apar Start_of_Run_Parity_918
  • 64291 06/23/01 16:30 R. Michaels re: "mysteries" of qasy
  • 64289 06/23/01 15:59 Binet Compton chicane problem
  • 64286 06/23/01 15:44 apar Start_of_Run_Parity_917
  • 64283 06/23/01 15:33 Seonho Choi Fast Feedback Off
  • 64282 06/23/01 15:28 Seonho Choi Automatic feedback for charge asymmetry
  • 64279 06/23/01 15:11 apar Start_of_Run_Parity_916
  • 64277 06/23/01 15:05 Binet huge counting rates in the compton photon detector
  • 64274 06/23/01 14:04 apar Start_of_Run_Parity_915
  • 64270 06/23/01 13:46 A. Deur Charge asymmetry screen was frozen
  • 64268 06/23/01 12:58 apar Start_of_Run_Parity_914
  • 64263 06/23/01 12:04 A. Deur out of sync run numbers
  • 64261 06/23/01 11:45 apar Start_of_Run_Parity_913
  • 64257 06/23/01 11:35 A. Deur fast feedback off
  • 64256 06/23/01 11:15 xfzhu , Karl EPR and NMR measurement
  • 64255 06/23/01 10:46 xfzhu ROC 14 tripped
  • 64253 06/23/01 10:32 apar Start_of_Run_Parity_912
  • 64248 06/23/01 08:54 apar Start_of_Run_Parity_911
  • 64245 06/23/01 08:12 X. Zheng Delta asymmetry discrepancy
  • 64243 06/23/01 08:02 Hong Xiang Shift Summary
  • 64240 06/23/01 07:35 apar Start_of_Run_Parity_910
  • 64237 06/23/01 07:18 X. Zheng x=0.6, asymmetry
  • 64236 06/23/01 06:52 X. Zheng x=0.6, longitudinal asymmetry @ target 180 deg
  • 64235 06/23/01 06:50 X. Zheng x=0.6, longitudinal asymmetry @ target 0 deg
  • 64234 06/23/01 06:47 X. Zheng x=0.6, transverse asymmetry
  • 64233 06/23/01 06:37 X. Zheng June 23 summary of x=0.6 electrons
  • 64230 06/23/01 06:22 apar Start_of_Run_Parity_909
  • 64227 06/23/01 06:14 X. Zheng Target - EPR - 36.0%
  • 64226 06/23/01 06:00 adaq Runsummary_Left_HRS
  • 64225 06/23/01 06:00 adaq Runsummary_Right_HRS
  • 64222 06/23/01 05:06 apar Start_of_Run_Parity_908
  • 64217 06/23/01 04:05 apar Start_of_Run_Parity_907
  • 64212 06/23/01 02:49 apar Start_of_Run_Parity_906
  • 64209 06/23/01 02:00 X. Zheng /K. Target - EPR 36.9%, NMR 36.5%, EPR 35.9%
  • 64206 06/23/01 01:39 apar Start_of_Run_Parity_905
  • 64203 06/23/01 00:09 Aniol shift summary
  • Friday

  • 64199 06/22/01 23:59 apar Start_of_Run_Parity_904
  • 64196 06/22/01 23:11 Armstrong charge asymmetry changes
  • 64193 06/22/01 22:55 apar Start_of_Run_Parity_903
  • 64190 06/22/01 22:11 Sulkosky Stray Run Numbers in Left Arm Analysis
  • 64189 06/22/01 22:02 Armstrong drifts in charge asymmetry
  • 64185 06/22/01 21:51 apar Start_of_Run_Parity_902
  • 64182 06/22/01 21:49 Armstrong charge asymmetry
  • 64181 06/22/01 21:15 xfzhu transverse gradient field reversed
  • 64180 06/22/01 21:00 Binet Compton Polarimetry shift summary
  • 64178 06/22/01 20:48 apar Start_of_Run_Parity_901
  • 64175 06/22/01 20:28 X. Zheng Target - EPR - 36.8%
  • 64174 06/22/01 20:23 Armstrong Charge asymmetry
  • 64173 06/22/01 20:16 Sulkosky Target-EPR Measurement
  • 64172 06/22/01 20:01 Aniol lasers off for epr
  • 64170 06/22/01 19:32 apar Start_of_Run_Parity_900
  • 64164 06/22/01 18:25 apar Start_of_Run_Parity_899
  • 64163 06/22/01 18:12 Sulkosky Target-He-3 position
  • 64162 06/22/01 18:11 Aniol beam back
  • 64157 06/22/01 18:05 apar Start_of_Run_Parity_898
  • 64154 06/22/01 17:39 Sulkosky Target-Moved to BeO
  • 64153 06/22/01 17:00 Armstrong start cosmic run
  • 64152 06/22/01 16:50 Aniol chopper slit
  • 64150 06/22/01 16:48 apar Start_of_Run_Parity_897
  • 64147 06/22/01 15:58 Ole Hansen Shift Summary
  • 64146 06/22/01 15:05 Wolfgang, X. Target: EPR
  • 64142 06/22/01 14:05 apar Start_of_Run_Parity_896
  • 64138 06/22/01 12:51 xfzhu CODA restarted for both arms
  • 64137 06/22/01 12:38 apar Start_of_Run_Parity_895
  • 64134 06/22/01 12:18 Wolfgang, B. Target: EPR, NMR, field rotation to 180 deg
  • 64127 06/22/01 10:07 apar Start_of_Run_Parity_894
  • 64121 06/22/01 08:50 apar Start_of_Run_Parity_893
  • 64117 06/22/01 08:07 kramerk Shift Summary
  • 64112 06/22/01 07:15 kramerk Target - NMR polarization 35.3%
  • 64108 06/22/01 06:54 X. Zheng Target - EPR 38.5%
  • 64106 06/22/01 06:00 adaq Runsummary_Left_HRS
  • 64105 06/22/01 06:00 adaq Runsummary_Right_HRS
  • 64103 06/22/01 05:49 apar Start_of_Run_Parity_891
  • 64098 06/22/01 04:47 apar Start_of_Run_Parity_890
  • 64092 06/22/01 03:30 apar Start_of_Run_Parity_889
  • 64089 06/22/01 02:49 kramerk Target - NMR Polarization 36.0%
  • 64088 06/22/01 02:43 X. Zheng Target - EPR, 39.6%
  • 64086 06/22/01 02:22 apar Start_of_Run_Parity_888
  • 64081 06/22/01 01:09 apar Start_of_Run_Parity_887
  • 64078 06/22/01 00:16 kramerk Beam is down for a few minutes
  • Thursday

  • 64076 06/21/01 23:54 K. McCormick Swing Summary June 21
  • 64075 06/21/01 23:54 K. McCormick Automatic charge feedback turned off for a second
  • 64072 06/21/01 23:41 apar Start_of_Run_Parity_886
  • 64071 06/21/01 22:37 kramerk Target - Changed Long. Coils to -2.0 Amps
  • 64070 06/21/01 22:34 X. Zheng Target - EPR - 38.9%
  • 64069 06/21/01 22:26 X. Zheng right arm shower/preshower update
  • 64068 06/21/01 22:01 kramerk Target - NMR polarization 36.0 %
  • 64067 06/21/01 20:43 K. McCormick Right Arm Shower/Preshower sparcely populated
  • 64066 06/21/01 20:36 J. Roche NMR measurement
  • 64060 06/21/01 19:41 apar Start_of_Run_Parity_885
  • 64059 06/21/01 19:40 Binet Compton Polarimetry Shift Summary
  • 64057 06/21/01 19:10 R. Michaels charge feedback enabled
  • 64054 06/21/01 18:38 apar Start_of_Run_Parity_884
  • 64048 06/21/01 17:37 apar Start_of_Run_Parity_883
  • 64044 06/21/01 17:31 X. Zheng, W. EPR - 38.96%
  • 64042 06/21/01 17:19 apar Start_of_Run_Parity_882
  • 64038 06/21/01 17:11 apar Start_of_Run_Parity_881
  • 64037 06/21/01 16:58 kramerk Target - NMR Polarization - 36.1 %
  • 64035 06/21/01 16:05 Sulkosky Shift Summary
  • 64033 06/21/01 14:58 R. Michaels Charge asymmetries past 24 hrs
  • 64027 06/21/01 13:48 apar Start_of_Run_Parity_880
  • 64022 06/21/01 12:43 apar Start_of_Run_Parity_879
  • 64020 06/21/01 11:51 Sulkosky Beam back
  • 64016 06/21/01 11:40 apar Start_of_Run_Parity_878
  • 64014 06/21/01 11:27 Sulkosky Beam off
  • 64012 06/21/01 11:09 Sulkosky Checked beam spot
  • 64009 06/21/01 11:01 apar Start_of_Run_Parity_877
  • 64004 06/21/01 10:52 apar Start_of_Run_Parity_876
  • 64001 06/21/01 10:30 Sulkosky Cosmic runs
  • 64000 06/21/01 09:49 Target EPR
  • 63999 06/21/01 09:47 X. Zheng Increase transvers gradient coil current
  • 63998 06/21/01 09:41 X. Zheng Target - EPR - 38.1%
  • 63995 06/21/01 08:58 apar Start_of_Run_Parity_875
  • 63988 06/21/01 07:58 Dutta Shift summary
  • 63986 06/21/01 07:37 xfzhu Reboot the Xterm for target control
  • 63983 06/21/01 06:00 adaq Runsummary_Left_HRS
  • 63982 06/21/01 06:00 adaq Runsummary_Right_HRS
  • 63979 06/21/01 05:56 apar Start_of_Run_Parity_874
  • 63977 06/21/01 05:33 xfzhu Target - EPR - 38.8%
  • 63973 06/21/01 04:51 apar Start_of_Run_Parity_873
  • 63968 06/21/01 03:47 apar Start_of_Run_Parity_872
  • 63963 06/21/01 02:45 apar Start_of_Run_Parity_871
  • 63960 06/21/01 02:04 X. Zheng June 20 summary of half waveplates/holding field
  • 63957 06/21/01 01:37 apar Start_of_Run_Parity_870
  • 63956 06/21/01 01:27 xfzhu EPR polarization extraction
  • 63952 06/21/01 00:52 X. Zheng Target - EPR - 39.6%
  • 63950 06/21/01 00:52 apar Start_of_Run_Parity_869
  • Wednesday

  • 63944 06/20/01 23:53 aniol shift summary
  • 63943 06/20/01 23:48 aniol beam wave plate change
  • 63940 06/20/01 23:44 apar Start_of_Run_Parity_868
  • 63936 06/20/01 22:56 apar Start_of_Run_Parity_867
  • 63931 06/20/01 21:41 apar Start_of_Run_Parity_866
  • 63924 06/20/01 20:28 apar Start_of_Run_Parity_865
  • 63920 06/20/01 20:17 jpchen fitted EPR result
  • 63919 06/20/01 19:53 jpchen EPR
  • 63917 06/20/01 19:17 apar Start_of_Run_Parity_864
  • 63912 06/20/01 19:02 apar Start_of_Run_Parity_863
  • 63909 06/20/01 18:22 jpchen target half-wave plates re-installed
  • 63907 06/20/01 17:55 apar Start_of_Run_Parity_862
  • 63901 06/20/01 17:44 apar Start_of_Run_Parity_861
  • 63900 06/20/01 16:35 xfzhu paw kumac for detector checking
  • 63898 06/20/01 16:09 jgao Day shift summary
  • 63894 06/20/01 15:53 X. Zheng Half waveplates test done
  • 63891 06/20/01 15:24 apar Start_of_Run_Parity_860
  • 63888 06/20/01 14:29 Wolfgang EPR
  • 63886 06/20/01 14:23 apar Start_of_Run_Parity_859
  • 63884 06/20/01 13:55 A. Deur delta asymmetry
  • 63879 06/20/01 13:37 apar Start_of_Run_Parity_858
  • 63878 06/20/01 13:27 A. Deur He3 pressure curve
  • 63874 06/20/01 12:23 apar Start_of_Run_Parity_857
  • 63870 06/20/01 12:01 R. Michaels parity DAQ turned on again
  • 63869 06/20/01 11:45 apar Start_of_Run_Parity_856
  • 63856 06/20/01 09:01 Wolfgang EPR
  • 63851 06/20/01 08:00 kramerk Shift Summary
  • 63841 06/20/01 06:00 adaq Runsummary_Left_HRS
  • 63840 06/20/01 06:00 adaq Runsummary_Right_HRS
  • 63832 06/20/01 04:30 X. Zheng summary of x=0.6 runs
  • 63830 06/20/01 03:42 kramerk Target - EPR measurement - 37.1%
  • 63825 06/20/01 01:28 kramerk Target - NMR Polarizations from 4pm and 8pm
  • 63824 06/20/01 01:10 kramerk Began Pol. 3He Running
  • 63823 06/20/01 01:10 kramerk Checked Beam Spot
  • Tuesday

  • 63813 06/19/01 23:41 S. Sirca Summary of evening shift
  • 63812 06/19/01 22:50 Dieterich MCC update (II)
  • 63811 06/19/01 22:47 The Target C Target: Addition of transverse gradient, EPR
  • 63809 06/19/01 22:06 apar Start_of_Run_Parity_855
  • 63807 06/19/01 21:39 apar Start_of_Run_Parity_854
  • 63803 06/19/01 21:36 apar Start_of_Run_Parity_853
  • 63801 06/19/01 21:05 Piotr and Wo NMR, EPR
  • 63800 06/19/01 20:42 dieterich MCC update
  • 63799 06/19/01 18:11 Wolfgang, X. NMR, EPR
  • 63796 06/19/01 17:37 S. Sirca Vacuum problem in N Linac
  • 63795 06/19/01 16:21 apar Start_of_Run_Parity_852
  • 63794 06/19/01 16:13 Sulkosky Day Shift Summary
  • 63792 06/19/01 15:32 X. Zheng half waveplates are at EEL lab now
  • 63791 06/19/01 15:14 E.Chudakov Shower (Right Arm) channel 79 problem
  • 63790 06/19/01 14:47 X. Zheng Remove half waveplates from laser hut
  • 63789 06/19/01 14:44 Sulkosky Beam-update
  • 63788 06/19/01 14:44 X. Zheng Rotate half waveplates back
  • 63787 06/19/01 14:19 R. Michaels modification to epics_end script, R-arm
  • 63786 06/19/01 11:26 Sulkosky Beam -off
  • 63785 06/19/01 10:59 Wolfgang EPR
  • 63784 06/19/01 10:51 Sulkosky Beam out for two hours
  • 63783 06/19/01 10:38 Sulkosky Beam off-Magnet Trips
  • 63776 06/19/01 09:48 kramerk Target - NMR Polarization 23.3%
  • 63775 06/19/01 09:32 Wolfgang NMR
  • 63771 06/19/01 08:32 X. Zheng Summary of field direction
  • 63764 06/19/01 08:00 jgao Owl shift summary
  • 63763 06/19/01 07:29 kramerk Target - EPR Polarization @ 4:45am - 34.4%
  • 63758 06/19/01 07:03 kramerk Target -NMR polarizations for 18 Jun
  • 63753 06/19/01 06:00 adaq Runsummary_Left_HRS
  • 63752 06/19/01 06:00 adaq Runsummary_Right_HRS
  • 63751 06/19/01 05:52 kramerk Target - NMR Polarization @ 5am - 31.4%
  • 63750 06/19/01 05:11 kramerk Target - Set Up in 180 degree position
  • 63747 06/19/01 04:57 kramerk Target - EPR measurement
  • 63728 06/19/01 00:10 Wolfgang fitted EPR result
  • Monday

  • 63727 06/18/01 23:56 S. Sirca Summary of evening shift
  • 63725 06/18/01 23:48 Wolfgang EPR
  • 63722 06/18/01 22:48 Eugene Moller results
  • 63720 06/18/01 20:59 S. Sirca Change in pumping strategy
  • 63718 06/18/01 20:52 X. Zheng, W. EPR measurement at 0 deg.
  • 63717 06/18/01 20:16 S. Sirca Quick update
  • 63713 06/18/01 19:33 apar Start_of_Run_Parity_851
  • 63712 06/18/01 19:21 apar Start_of_Run_Parity_850
  • 63711 06/18/01 19:19 apar Start_of_Run_Parity_849
  • 63708 06/18/01 19:02 S. Sirca Temporary change of run-plan
  • 63699 06/18/01 18:08 X. Zheng Laser Work during Maintaince
  • 63698 06/18/01 16:06 apar Start_of_Run_Parity_848
  • 63697 06/18/01 16:01 apar Start_of_Run_Parity_847
  • 63695 06/18/01 16:00 LeRose Shift Summary
  • 63694 06/18/01 15:54 apar Start_of_Run_Parity_846
  • 63693 06/18/01 15:49 apar Start_of_Run_Parity_845
  • 63692 06/18/01 15:42 apar Start_of_Run_Parity_844
  • 63691 06/18/01 15:36 apar Start_of_Run_Parity_843
  • 63690 06/18/01 15:31 apar Start_of_Run_Parity_842
  • 63689 06/18/01 15:29 apar Start_of_Run_Parity_841
  • 63688 06/18/01 15:22 apar Start_of_Run_Parity_840
  • 63687 06/18/01 15:14 apar Start_of_Run_Parity_839
  • 63686 06/18/01 14:57 apar Start_of_Run_Parity_838
  • 63685 06/18/01 14:55 apar Start_of_Run_Parity_837
  • 63683 06/18/01 14:48 apar Start_of_Run_Parity_835
  • 63682 06/18/01 14:37 apar Start_of_Run_Parity_834
  • 63681 06/18/01 14:15 apar Start_of_Run_Parity_833
  • 63680 06/18/01 13:58 Wolfgang Last EPR result
  • 63678 06/18/01 13:46 Wolfgang current in grad. coil changed (in the morning)
  • 63677 06/18/01 13:41 LeRose Beam Off
  • 63676 06/18/01 13:00 LeRose Beam Off
  • 63674 06/18/01 12:39 LeRose No Beam
  • 63665 06/18/01 10:39 Wolfgang EPR, NMR
  • 63663 06/18/01 09:55 LeRose Target 180 deg NOT 0
  • 63662 06/18/01 09:43 LeRose Target 180 deg NOT 0
  • 63399 06/18/01 09:03 apar Start_of_Run_Parity_798
  • 63656 06/18/01 08:02 E.Chudakov Owl shift summary
  • 63651 06/18/01 07:22 kramerk Target - EPR Measurement 34.9%
  • 63650 06/18/01 07:12 E.Chudakov HV tripped along with the alarm IOC
  • 63646 06/18/01 06:41 X. Zheng correction
  • 63386 06/18/01 06:39 apar Start_of_Run_Parity_796
  • 63641 06/18/01 06:10 X. Zheng Summary of electrons, x=0.6, longitudinal - Settings I & II
  • 63639 06/18/01 06:00 adaq Runsummary_Left_HRS
  • 63638 06/18/01 06:00 adaq Runsummary_Right_HRS
  • 63637 06/18/01 05:12 E.Chudakov IOC for HV alarm needs rebooting
  • 63628 06/18/01 02:30 kramerk Target - Gradient Coils Current increased to -6.0Amps
  • 63627 06/18/01 02:29 kramerk Target - EPR Polarization -34.6 %
  • 63618 06/18/01 00:11 A. Powell summary of swing shift
  • Sunday

  • 63609 06/17/01 22:14 vsulkosky Target-EPR polarization 35.8 %@17:50
  • 63608 06/17/01 22:06 vsulkosky Gradient Field at -4.04 V
  • 63607 06/17/01 22:04 jpchen Chnaged gradient field to -4 V
  • 63606 06/17/01 22:03 vsulkosky Target-EPR 35.3%
  • 63590 06/17/01 18:38 Wolfgang EPR
  • 63593 06/17/01 18:32 V.Sulkosky Half-waveplate In and Target at Pol. 3He
  • 63592 06/17/01 18:27 V.Sulkosky Target-EPR 38 %
  • 63591 06/17/01 18:21 kramerk Target - Kevin's Polarimetry Plan
  • 63587 06/17/01 18:14 kramerk Target - Field rotated to 180
  • 63586 06/17/01 18:13 Antonin Vach DAQ runs DECOUPLED , parity DAQ pbs
  • 63585 06/17/01 18:06 apar Start_of_Run_Parity_831
  • 63584 06/17/01 18:04 apar Start_of_Run_Parity_830
  • 63326 06/17/01 17:47 apar Start_of_Run_Parity_789
  • 63579 06/17/01 17:41 apar Start_of_Run_Parity_829
  • 63578 06/17/01 17:14 apar Start_of_Run_Parity_828
  • 63577 06/17/01 17:13 apar Start_of_Run_Parity_827
  • 63576 06/17/01 16:58 apar Start_of_Run_Parity_826
  • 63572 06/17/01 15:58 S. Sirca Summary of day shift
  • 63568 06/17/01 15:45 apar Start_of_Run_Parity_825
  • 63564 06/17/01 14:44 apar Start_of_Run_Parity_824
  • 63560 06/17/01 14:01 Armstrong Beam Charge Asymmetry Summary
  • 63559 06/17/01 13:54 Armstrong beam charge asym - summary
  • 63556 06/17/01 13:40 apar Start_of_Run_Parity_823
  • 63554 06/17/01 13:36 S. Sirca Beam is back
  • 63553 06/17/01 13:19 S. Sirca Beam off, problems with trim lock
  • 63550 06/17/01 13:15 apar Start_of_Run_Parity_822
  • 63546 06/17/01 12:28 R. Michaels chg asy runs 20599 and 616
  • 63544 06/17/01 12:11 apar Start_of_Run_Parity_821
  • 63539 06/17/01 11:30 X. Zheng Target - EPR 37.20%
  • 63538 06/17/01 11:23 S. Sirca EPR measurement and new wave-plate position
  • 63536 06/17/01 11:07 apar Start_of_Run_Parity_820
  • 63533 06/17/01 10:41 S. Sirca Slow DAQ start/stop on Right HRS
  • 63530 06/17/01 10:38 apar Start_of_Run_Parity_819
  • 63525 06/17/01 09:34 apar Start_of_Run_Parity_818
  • 63522 06/17/01 08:45 S. Sirca Problems with charge asymmetries
  • 63520 06/17/01 08:31 apar Start_of_Run_Parity_817
  • 63514 06/17/01 07:37 kramerk Target-NMR Polarization 34.2%
  • 63512 06/17/01 07:28 X. Zheng Target - EPR 37.3%
  • 63510 06/17/01 07:27 apar Start_of_Run_Parity_816
  • 63504 06/17/01 07:18 apar Start_of_Run_Parity_815
  • 63498 06/17/01 07:01 apar Start_of_Run_Parity_814
  • 63495 06/17/01 06:18 Adaq Target - NMR Polarization @ 9:31pm = 34.2%
  • 63492 06/17/01 06:04 apar Start_of_Run_Parity_813
  • 63489 06/17/01 06:00 adaq Runsummary_Left_HRS
  • 63488 06/17/01 06:00 adaq Runsummary_Right_HRS
  • 63226 06/17/01 05:02 apar Start_of_Run_Parity_773
  • 63482 06/17/01 04:49 X. Zheng Summary of electrons, x=0.6, transverse
  • 63478 06/17/01 04:07 apar Start_of_Run_Parity_811
  • 63473 06/17/01 03:08 apar Start_of_Run_Parity_810
  • 63471 06/17/01 02:36 X. Zheng Target - EPR 37.7%
  • 63470 06/17/01 02:29 Z. Chai Charge Asymmetry feedback control
  • 63466 06/17/01 02:07 apar Start_of_Run_Parity_809
  • 63457 06/17/01 01:12 apar Start_of_Run_Parity_807
  • 63452 06/17/01 00:52 apar Start_of_Run_Parity_806
  • 63449 06/17/01 00:23 Nilanga Incorrect halog entry for runs 1586 and 20586 corrected
  • 63448 06/17/01 00:07 S. Sirca Erratum
  • Saturday

  • 63446 06/16/01 23:59 S. Sirca Summary of evening shift
  • 63443 06/16/01 23:36 apar Start_of_Run_Parity_805
  • 63441 06/16/01 22:35 Seonho Choi MCC, beam trip, chopping slit
  • 63439 06/16/01 21:42 W. Korsch & EPR, NMR
  • 63436 06/16/01 21:36 apar Start_of_Run_Parity_804
  • 63431 06/16/01 20:11 apar Start_of_Run_Parity_803
  • 63422 06/16/01 18:06 apar Start_of_Run_Parity_801
  • 63416 06/16/01 17:59 apar Start_of_Run_Parity_800
  • 63412 06/16/01 17:00 X. Zhu, Wolf EPR -target
  • 63408 06/16/01 16:56 apar Start_of_Run_Parity_799
  • 63407 06/16/01 16:52 Seonho Choi EPR measurement
  • 63404 06/16/01 15:16 Armstrong beam charge asymmetry
  • 63403 06/16/01 15:12 A. Deur Injector half wave plate in
  • 63399 06/16/01 15:09 apar Start_of_Run_Parity_798
  • 63395 06/16/01 14:11 apar Start_of_Run_Parity_797
  • 63392 06/16/01 13:36 Armstrong back to 10 muA
  • 63390 06/16/01 13:17 X. Zheng How many electrons do we have now?
  • 63389 06/16/01 12:51 Armstrong beam charge asym.- new value
  • 63386 06/16/01 12:49 apar Start_of_Run_Parity_796
  • 63383 06/16/01 12:29 X. Zheng Target - EPR with gradient coil test
  • 63382 06/16/01 12:19 kramerk Target - NMR Polarization 36.15
  • 63381 06/16/01 12:16 A. Deur Half wave plates status.
  • 63380 06/16/01 12:01 A. Deur Total number of electron for transverse data taking
  • 63377 06/16/01 11:56 apar Start_of_Run_Parity_795
  • 63628 06/16/01 11:50 kramerk Target - Gradient Coils Current increased to -6.0Amps
  • 63370 06/16/01 11:45 A. Deur Target is on longitudinal mode
  • 63369 06/16/01 11:37 kramerk Target - Rebooted Hall Computer
  • 63367 06/16/01 10:50 A. Deur Hall A in control access.
  • 63366 06/16/01 10:47 A. Deur Injector in control accecss
  • 63362 06/16/01 10:15 A. Deur beam current
  • 63101 06/16/01 10:15 apar Start_of_Run_Parity_755
  • 63360 06/16/01 10:14 apar Start_of_Run_Parity_793
  • 63356 06/16/01 10:07 Armstron recent beam charge asymmetries
  • 63355 06/16/01 09:32 A. Deur Beam current
  • 63351 06/16/01 09:07 X. Zheng Target - EPR
  • 63350 06/16/01 09:06 Armstrong Go to NMR and EPR
  • 63349 06/16/01 08:52 kramerk Target - NMR Polarization 36.59%
  • 63346 06/16/01 08:19 Armstrong FastFeedback
  • 63345 06/16/01 08:19 X.Zheng shift summary
  • 63339 06/16/01 07:20 apar Start_of_Run_Parity_791
  • 63079 06/16/01 07:16 Armstrong FastFeedback adjustment
  • 63337 06/16/01 06:47 kramerk Target - Commentary on NMR Gradient Fits
  • 63335 06/16/01 06:22 apar Start_of_Run_Parity_790
  • 63331 06/16/01 06:00 adaq Runsummary_Left_HRS
  • 63330 06/16/01 06:00 adaq Runsummary_Right_HRS
  • 63329 06/16/01 05:52 kramerk Target - NMR Fit with Gradient
  • 63326 06/16/01 05:23 apar Start_of_Run_Parity_789
  • 63323 06/16/01 04:45 Camsonne A. Reference cell
  • 63322 06/16/01 04:23 Camsonne A. Target 270 degres
  • 63321 06/16/01 04:16 X. Zheng Target - EPR 37.3%
  • 63576 06/16/01 04:05 apar Start_of_Run_Parity_826
  • 63051 06/16/01 02:02 Eugene Moller results
  • 63310 06/16/01 01:18 Camsonne A. Beam back
  • 63309 06/16/01 00:44 Camsonne A. Accelerator RF problem
  • Friday

  • 63307 06/15/01 23:53 ddutta Shift Summary
  • 63303 06/15/01 23:31 apar Start_of_Run_Parity_786
  • 63302 06/15/01 22:28 Z. Chai Target change to Reference Cell
  • 63301 06/15/01 21:38 Wolfgang, X. NMR, EPR, NMR
  • 63298 06/15/01 19:29 apar Start_of_Run_Parity_785
  • 63291 06/15/01 19:19 apar Start_of_Run_Parity_784
  • 63288 06/15/01 17:49 apar Start_of_Run_Parity_783
  • 63023 06/15/01 17:48 jpchen Swing Shift summary
  • 63282 06/15/01 16:47 apar Start_of_Run_Parity_782
  • 63279 06/15/01 16:28 Wolfgang EPR
  • 63014 06/15/01 16:10 apar Start_of_Run_Parity_744
  • 63277 06/15/01 16:01 Saha Shift Summary
  • 63274 06/15/01 15:48 apar Start_of_Run_Parity_781
  • 63525 06/15/01 14:44 apar Start_of_Run_Parity_818
  • 63269 06/15/01 14:44 apar Start_of_Run_Parity_780
  • 63265 06/15/01 14:08 kramerk Target - 10:53am NMR Polarization
  • 63520 06/15/01 13:40 apar Start_of_Run_Parity_817
  • 63258 06/15/01 12:36 apar Start_of_Run_Parity_778
  • 63253 06/15/01 12:32 apar Start_of_Run_Parity_777
  • 63249 06/15/01 12:10 Wolfgang Kor EPR Measurement
  • 63248 06/15/01 11:47 Zolnierczuk TARGET NMR/EPR
  • 63247 06/15/01 11:27 Adev raster at 5 uAmps
  • 63241 06/15/01 11:20 Binet Raster check
  • 63240 06/15/01 11:19 Binet Raster check
  • 63238 06/15/01 11:15 apar Start_of_Run_Parity_775
  • 63233 06/15/01 09:42 apar Start_of_Run_Parity_774
  • 62958 06/15/01 09:41 adaq Runsummary_Left_HRS
  • 63227 06/15/01 08:40 Binet Raster shape while harp scan
  • 63226 06/15/01 08:38 apar Start_of_Run_Parity_773
  • 63478 06/15/01 08:31 apar Start_of_Run_Parity_811
  • 63220 06/15/01 08:31 apar Start_of_Run_Parity_772
  • 63214 06/15/01 07:56 R. Michaels owl shift
  • 63213 06/15/01 07:52 X. Zheng Target - Helmholtz coil current
  • 63211 06/15/01 07:25 apar Start_of_Run_Parity_771
  • 63208 06/15/01 06:44 kramerk Target - Holding Field getting smaller
  • 63207 06/15/01 06:42 kramerk Target - Gradient Coil NMR Test
  • 63205 06/15/01 06:13 apar Start_of_Run_Parity_770
  • 63203 06/15/01 06:11 Camsonne A. Target He3 Transverse
  • 62932 06/15/01 06:00 kramerk Target - NMR Measurement
  • 63200 06/15/01 05:17 R. Michaels rotatable waveplate adjustment
  • 62925 06/15/01 05:04 R. Michaels beam back, HARP, OTR, spot size
  • 63196 06/15/01 05:04 R. Michaels poor beam quality ?
  • 63195 06/15/01 04:41 R. Michaels slope for feedback
  • 63194 06/15/01 04:28 Adaq puzzle - forgot description
  • 63193 06/15/01 04:25 X.Zheng Puzzle for Target group
  • 63192 06/15/01 04:02 R. Michaels correction: field 270 degrees
  • 63181 06/15/01 03:30 R. Michaels run 20567 charge asy.
  • 63179 06/15/01 02:47 Camsonne A. Pulsed beam
  • 63176 06/15/01 02:30 apar Start_of_Run_Parity_767
  • 63170 06/15/01 02:15 apar Start_of_Run_Parity_766
  • 63164 06/15/01 02:00 X. Zheng Target - EPR etc.
  • 63416 06/15/01 01:52 apar Start_of_Run_Parity_800
  • 63152 06/15/01 01:37 apar Start_of_Run_Parity_763
  • 63144 06/15/01 00:49 apar Start_of_Run_Parity_761
  • 62867 06/15/01 00:44 adaq Runsummary_Left_HRS
  • 63141 06/15/01 00:40 Camsonne A. Hall B test
  • 63139 06/15/01 00:18 A. Deur N2 Pressure Curve
  • 63138 06/15/01 00:17 xfzhu NMR measurement
  • 63137 06/15/01 00:14 A. Deur Elastic cross section
  • 63136 06/15/01 00:02 ddutta Shift Summary
  • Thursday

  • 63133 06/14/01 23:58 A. Deur Elastic asymmetry
  • 63130 06/14/01 23:54 apar Start_of_Run_Parity_760
  • 63125 06/14/01 22:51 apar Start_of_Run_Parity_759
  • 63122 06/14/01 22:11 Wolfgang EPR, what else?
  • 63377 06/14/01 22:01 apar Start_of_Run_Parity_795
  • 63114 06/14/01 20:48 apar Start_of_Run_Parity_757
  • 62839 06/14/01 20:44 apar Start_of_Run_Parity_725
  • 63366 06/14/01 19:37 A. Deur Injector in control accecss
  • 63104 06/14/01 18:28 jpchen possible mzing and gradient field on now
  • 63101 06/14/01 18:22 apar Start_of_Run_Parity_755
  • 63100 06/14/01 18:22 Z. Chai New target configuration
  • 63093 06/14/01 18:11 apar Start_of_Run_Parity_754
  • 63090 06/14/01 18:05 apar Start_of_Run_Parity_753
  • 63089 06/14/01 18:05 jpchen lasers
  • 63346 06/14/01 18:04 Armstrong FastFeedback
  • 63082 06/14/01 16:16 X. Zheng Laser power and waveplates check
  • 63081 06/14/01 16:09 Zolnierczuk day shift summary
  • 63079 06/14/01 14:36 Armstrong FastFeedback adjustment
  • 63078 06/14/01 14:31 Armstrong Rotate holding field to 270 deg.
  • 62806 06/14/01 14:02 apar Start_of_Run_Parity_719
  • 63076 06/14/01 13:50 Armstrong EPR measurement
  • 63075 06/14/01 13:15 Armstrong Beam off - NMR measurement
  • 63072 06/14/01 12:24 apar Start_of_Run_Parity_751
  • 63070 06/14/01 12:24 Armstrong charge asymmetry
  • 63064 06/14/01 11:10 apar Start_of_Run_Parity_750
  • 63061 06/14/01 11:00 Armstrong production runs restarted
  • 63057 06/14/01 10:55 apar Start_of_Run_Parity_749
  • 63056 06/14/01 10:41 Binet raster spot
  • 63053 06/14/01 10:35 apar Start_of_Run_Parity_748
  • 63052 06/14/01 08:47 Armstrong EPR measurement
  • 63051 06/14/01 08:22 Eugene Moller results
  • 63050 06/14/01 08:22 R. Michaels owl shift
  • 63049 06/14/01 08:19 Armstrong move to BeO target
  • 63048 06/14/01 08:17 Eugene Moller results
  • 63046 06/14/01 07:09 R. Michaels Moeller chg asy
  • 63045 06/14/01 06:28 R. Michaels relative sign calibration
  • 63042 06/14/01 06:00 adaq Runsummary_Left_HRS
  • 63037 06/14/01 05:59 apar Start_of_Run_Parity_747
  • 63036 06/14/01 05:48 X. Zheng x=0.6, longitudinal asymmetry
  • 63033 06/14/01 05:08 apar Start_of_Run_Parity_746
  • 63030 06/14/01 03:46 X. Zheng delta, transverse asymmetry
  • 63029 06/14/01 03:37 Camsonne A. Target NMR 32.42%
  • 63023 06/14/01 00:08 jpchen Swing Shift summary
  • 62746 06/14/01 00:02 adaq Runsummary_Right_HRS
  • Wednesday

  • 63019 06/13/01 23:38 jpchen Compton beam vertical turning
  • 63016 06/13/01 23:08 jpchen EPR and 3 lasers now for longitudinal pumping
  • 63015 06/13/01 22:24 jpchen beam trips reduced
  • 63014 06/13/01 22:04 apar Start_of_Run_Parity_744
  • 63009 06/13/01 21:44 apar Start_of_Run_Parity_743
  • 63525 06/13/01 21:32 apar Start_of_Run_Parity_818
  • 63265 06/13/01 21:25 kramerk Target - 10:53am NMR Polarization
  • 62999 06/13/01 21:05 jpchen correction to status
  • 62998 06/13/01 21:03 R. Michaels slightly better qasy analysis
  • 62997 06/13/01 21:00 jpchen status
  • 62996 06/13/01 20:54 R. Michaels restored Parity DAQ
  • 62995 06/13/01 20:44 apar Start_of_Run_Parity_740
  • 62994 06/13/01 20:43 apar Start_of_Run_Parity_742
  • 62986 06/13/01 19:32 V. Sulkosky RTD GUI Alarm
  • 62985 06/13/01 19:15 V. Sulkosky NMR Polarization Extraction
  • 62984 06/13/01 18:38 jpchen HARP scan
  • 62983 06/13/01 18:35 V. Sulkosky NMR and EPR Measurements
  • 62982 06/13/01 18:25 jpchen charge asymmetry check
  • 62981 06/13/01 18:22 jpchen Happex DAQ not working now
  • 62980 06/13/01 18:19 jpchen activities since 4:30 pm.
  • 62975 06/13/01 17:00 jpchen Beam centering on BeO
  • 62974 06/13/01 16:49 jpchen Caution on BeO target
  • 62973 06/13/01 16:40 jpchen target position and motion FSD
  • 62972 06/13/01 16:20 jpchen Target vertical position encode
  • 62971 06/13/01 16:04 gilman shift summary
  • 62969 06/13/01 15:50 X. Zheng Gas cherenkov PID performance
  • 62967 06/13/01 14:46 gilman rad water leak, no beam until ~ 4pm
  • 62965 06/13/01 12:36 jpchen Target motion
  • 62964 06/13/01 11:06 gilman current status
  • 62963 06/13/01 07:57 K. McCormick Wed. OWL Shift Summary
  • 62961 06/13/01 07:28 R. Michaels auto FB working now
  • 62960 06/13/01 07:05 kramerk Several Target Lifting problems
  • 62959 06/13/01 06:17 R. Michaels Chg Asy, auto FB not working
  • 62958 06/13/01 06:00 adaq Runsummary_Left_HRS
  • 62956 06/13/01 05:53 kramerk Target - NMR Measurement
  • 62952 06/13/01 04:18 apar Start_of_Run_Parity_737
  • 62949 06/13/01 03:48 xfzhu Right arm shower calibration
  • 62946 06/13/01 03:16 apar Start_of_Run_Parity_736
  • 62666 06/13/01 03:02 Armstrong charge asymmetry confusion
  • 62942 06/13/01 02:49 R. Michaels multiple peaks in scint TDCs (no problem)
  • 63208 06/13/01 02:25 kramerk Target - Holding Field getting smaller
  • 63207 06/13/01 02:24 kramerk Target - Gradient Coil NMR Test
  • 62935 06/13/01 02:18 Nilanga Summary of target activities during the maintanance data
  • 62934 06/13/01 01:51 Nilanga espace_kumac restored
  • 62933 06/13/01 01:40 X. Zheng Target - EPR
  • 62932 06/13/01 01:34 kramerk Target - NMR Measurement
  • 62930 06/13/01 01:26 apar Start_of_Run_Parity_734
  • 62928 06/13/01 01:16 X. Zheng hbook files
  • 62927 06/13/01 01:11 K. McCormick Beam position on target
  • 62926 06/13/01 00:56 K. McCormick Harp Scan with Raster On
  • 62925 06/13/01 00:55 R. Michaels beam back, HARP, OTR, spot size
  • 62922 06/13/01 00:51 apar Start_of_Run_Parity_733
  • 62919 06/13/01 00:42 X. Zheng espace_kumac directory
  • 62918 06/13/01 00:41 K. McCormick Harp Scan
  • Tuesday

  • 62913 06/12/01 23:50 S. Sirca Evening shift summary
  • 63179 06/12/01 21:09 Camsonne A. Pulsed beam
  • 62906 06/12/01 21:08 apar Start_of_Run_Parity_731
  • 62902 06/12/01 17:02 apar Start_of_Run_Parity_730
  • 62899 06/12/01 17:00 atrig Trigger-Download
  • 62897 06/12/01 16:47 atrig Trigger-Download
  • 62895 06/12/01 15:48 Zolnierczuk TARGET: NMR
  • 62894 06/12/01 12:31 X. Zheng x=0.6, transverse asymmetry
  • 62893 06/12/01 12:29 X. Zheng Delta asymmetry
  • 62891 06/12/01 10:59 Zolnierczuk TARGET NMR/EPR
  • 62886 06/12/01 06:00 adaq Runsummary_Right_HRS
  • Monday

  • 62883 06/11/01 20:02 xfzhu problem with laser #4
  • 62882 06/11/01 19:47 X. Zheng Target - EPR
  • 62881 06/11/01 19:34 xfzhu NMR polarization
  • 62880 06/11/01 18:18 K. McCormick Vertical Target Lifting
  • 62878 06/11/01 17:12 busato pion rejector calibration updated
  • 62877 06/11/01 17:10 jpchen Helium flow starts again
  • 62875 06/11/01 15:02 jpchen/xfzhu laser wavelength check and adjustment
  • 62873 06/11/01 11:59 apar Start_of_Run_Parity_729
  • 62871 06/11/01 11:43 R. Michaels L-arm T1 scaling fixed
  • 62868 06/11/01 07:00 X. Zheng Target - EPR - photodiode
  • 62867 06/11/01 06:00 adaq Runsummary_Left_HRS
  • 62866 06/11/01 06:00 adaq Runsummary_Right_HRS
  • 62865 06/11/01 01:14 xfzhu NMR polarization
  • 62864 06/11/01 00:52 X. Zheng Target - EPR
  • 62863 06/11/01 00:11 strauch Swing shift summary
  • Sunday

  • 62856 06/10/01 23:04 apar Start_of_Run_Parity_728
  • 62848 06/10/01 21:02 xfzhu NMR polarization: 40.36%
  • 62847 06/10/01 21:01 R. Michaels chkfeedback script
  • 62570 06/10/01 20:52 apar Parity_Start_of_Run_681
  • 62844 06/10/01 20:51 apar Start_of_Run_Parity_726
  • 62839 06/10/01 19:38 apar Start_of_Run_Parity_725
  • 62830 06/10/01 17:36 apar Start_of_Run_Parity_723
  • 62825 06/10/01 17:25 apar Start_of_Run_Parity_722
  • 63089 06/10/01 15:31 jpchen lasers
  • 62816 06/10/01 15:14 Camsonne A. Beam back
  • 62812 06/10/01 15:08 apar Start_of_Run_Parity_720
  • 62811 06/10/01 14:44 Camsonne A. Target He3 Polarized
  • 62810 06/10/01 14:37 Armstrong half wave plate out
  • 62806 06/10/01 14:28 apar Start_of_Run_Parity_719
  • 63076 06/10/01 14:26 Armstrong EPR measurement
  • 62804 06/10/01 14:05 Camsonne A. Target Beo - Beam coming back
  • 62803 06/10/01 14:04 Camsonne A. Target NMR
  • 62802 06/10/01 12:56 Armstrong more on charge asym
  • 62801 06/10/01 12:27 Camsonne A. Beam down
  • 63070 06/10/01 12:20 Armstrong charge asymmetry
  • 62794 06/10/01 11:33 apar Start_of_Run_Parity_718
  • 62793 06/10/01 11:31 Camsonne A. Beam back
  • 62792 06/10/01 11:15 Camsonne A. Target moved to polarized He3
  • 62791 06/10/01 11:10 Armstrong Charge Asymmetry
  • 62788 06/10/01 10:55 Armstrong Fast Feedback Checks
  • 62784 06/10/01 10:46 apar Start_of_Run_Parity_717
  • 62781 06/10/01 10:44 Camsonne A. Target moved to Beo
  • 62779 06/10/01 10:40 apar Start_of_Run_Parity_716
  • 62776 06/10/01 10:26 Camsonne A. Target Laser intensity Transverse mode
  • 62775 06/10/01 10:12 Armstrong Fast Feedback
  • 62774 06/10/01 10:08 Armstrong Charge asym feedback fixed?
  • 62773 06/10/01 09:36 apar Start_of_Run_Parity_715
  • 62766 06/10/01 09:12 apar Start_of_Run_Parity_714
  • 62368 06/10/01 08:00 xfzhu right arm shower/preshower calibration
  • 62756 06/10/01 07:07 R. Michaels re: strange left VDC tdc
  • 62755 06/10/01 07:02 X. Zheng Target - NMR
  • 62754 06/10/01 06:42 R. Michaels charge asymmetry feedback problems
  • 62752 06/10/01 06:39 apar Start_of_Run_Parity_712
  • 62746 06/10/01 06:00 adaq Runsummary_Right_HRS
  • 62742 06/10/01 05:16 apar Start_of_Run_Parity_711
  • 62740 06/10/01 05:06 Lingyan strange left vdc_tdc
  • 62736 06/10/01 04:06 apar Start_of_Run_Parity_710
  • 62735 06/10/01 03:24 R. Michaels beam position, showing dwell
  • 62731 06/10/01 03:00 apar Start_of_Run_Parity_709
  • 62727 06/10/01 02:17 R. Michaels beam back, production start, no FB
  • 62726 06/10/01 02:11 X. Zheng Target - last EPR
  • 62723 06/10/01 01:51 apar Start_of_Run_Parity_708
  • 62997 06/10/01 01:40 jpchen status
  • 62718 06/10/01 01:08 apar Start_of_Run_Parity_707
  • 62714 06/10/01 00:36 xfzhu NMR polarization
  • 62713 06/10/01 00:18 J. Roche shift summary
  • Saturday

  • 62711 06/09/01 23:28 apar Start_of_Run_Parity_706
  • 62705 06/09/01 21:57 apar Start_of_Run_Parity_705
  • 62700 06/09/01 21:19 apar Start_of_Run_Parity_704
  • 62695 06/09/01 20:09 apar Start_of_Run_Parity_703
  • 62693 06/09/01 19:30 xfzhu Transverse lasers used for 270 degree's setting
  • 62691 06/09/01 19:02 busato pion rejector calibration
  • 62689 06/09/01 18:34 xfzhu Holding field changed to 270 degree
  • 62687 06/09/01 18:32 apar Start_of_Run_Parity_702
  • 62685 06/09/01 18:19 busato pion rejector calibration
  • 62958 06/09/01 18:06 adaq Runsummary_Left_HRS
  • 62681 06/09/01 18:03 apar Start_of_Run_Parity_701
  • 62677 06/09/01 17:20 apar Start_of_Run_Parity_700
  • 62673 06/09/01 16:42 xfzhu NMR polarization
  • 62672 06/09/01 16:32 Nilanga 12C peaks at E_beam=5.7 GeV
  • 62669 06/09/01 15:42 Zolnierczuk run summary script
  • 62668 06/09/01 15:29 adaq Runsummary_Left_HRS
  • 62666 06/09/01 14:59 Armstrong charge asymmetry confusion
  • 62664 06/09/01 14:22 Zolnierczuk Parity DAQ: halog entry generation (fixed)
  • 62663 06/09/01 14:10 finn harp scan
  • 62661 06/09/01 14:04 apar Parity_Start_of_Run_692
  • 62659 06/09/01 14:01 Zolnierczuk TARGET: confusion
  • 62657 06/09/01 12:51 Zolnierczuk TARGET: NMR
  • 62654 06/09/01 12:12 Eugene Moller results
  • 62652 06/09/01 08:56 Armstrong go to Moller
  • 62651 06/09/01 08:41 Armstrong spot check
  • 62650 06/09/01 08:39 X. Zheng Delta asymmetry
  • 62648 06/09/01 08:30 R. Michaels OTR was not valid
  • 62646 06/09/01 08:22 finn HARP SCAN
  • 62640 06/09/01 07:41 R. Michaels BeO, no raster, HARP.
  • 62638 06/09/01 07:29 apar Parity_Start_of_Run_691
  • 62355 06/09/01 07:26 J.P./Piotr TARGET: He3 reference
  • 62633 06/09/01 05:38 X. Zheng Target - EPR
  • 62632 06/09/01 05:34 R. Michaels tune status, eP ion chambers
  • 62631 06/09/01 02:51 R. Michaels unplug daq synch. await 5 pass. update
  • 62627 06/09/01 01:34 Kelly file split on 1491
  • 62902 06/09/01 01:31 apar Start_of_Run_Parity_730
  • 62623 06/09/01 01:25 Kelly beam position drift
  • 62620 06/09/01 00:47 apar Parity_Start_of_Run_690
  • Friday

  • 62611 06/08/01 23:58 J. Roche Shift summary
  • 62609 06/08/01 23:52 apar Parity_Start_of_Run_688
  • 62216 06/08/01 23:32 apar Parity_Start_of_Run_625
  • 62877 06/08/01 23:31 jpchen Helium flow starts again
  • 62597 06/08/01 23:10 apar Parity_Start_of_Run_686
  • 62592 06/08/01 22:49 apar Parity_Start_of_Run_685
  • 62866 06/08/01 21:35 adaq Runsummary_Right_HRS
  • 62588 06/08/01 21:23 apar Parity_Start_of_Run_684
  • 62582 06/08/01 20:52 apar Parity_Start_of_Run_683
  • 62579 06/08/01 20:30 apar Parity_Start_of_Run_682
  • 62574 06/08/01 20:14 Zolnierczuk RUN SUMMARY TABLES
  • 62573 06/08/01 20:09 zolnie Left
  • 62572 06/08/01 20:09 zolnie Right
  • 62570 06/08/01 20:09 apar Parity_Start_of_Run_681
  • 62563 06/08/01 19:45 apar Parity_Start_of_Run_680
  • 62830 06/08/01 19:18 apar Start_of_Run_Parity_723
  • 62554 06/08/01 19:09 kramerk Target - Holding Field changed in the 180 degree Config.
  • 62552 06/08/01 18:56 apar Parity_Start_of_Run_678
  • 62548 06/08/01 18:20 Antonin Parity runs for feedback
  • 62547 06/08/01 18:16 X. Zheng Target EPR
  • 62545 06/08/01 18:10 apar Parity_Start_of_Run_677
  • 62543 06/08/01 17:49 apar Parity_Start_of_Run_676
  • 62542 06/08/01 17:36 apar Parity_Start_of_Run_675
  • 62541 06/08/01 17:26 apar Parity_Start_of_Run_674
  • 62816 06/08/01 17:23 Camsonne A. Beam back
  • 62538 06/08/01 17:17 apar Parity_Start_of_Run_673
  • 62537 06/08/01 17:04 apar Parity_Start_of_Run_672
  • 62536 06/08/01 16:57 apar Parity_Start_of_Run_671
  • 62810 06/08/01 16:45 Armstrong half wave plate out
  • 62533 06/08/01 16:26 Antonin Parity DAQ analyzer pictures !
  • 62248 06/08/01 16:22 Antonin Charge Asymmetry runs for feedbacks
  • 62529 06/08/01 16:21 Antonin Parity DAQ analyser is working fine !!
  • 62522 06/08/01 15:20 apar Parity_Start_of_Run_669
  • 62517 06/08/01 14:57 apar Parity_Start_of_Run_668
  • 62791 06/08/01 14:57 Armstrong Charge Asymmetry
  • 62513 06/08/01 14:37 Zolnierczuk Charge asymmetry
  • 62119 06/08/01 14:35 Ole Hansen Reposition beam at BPMA to minimize beam trajectory angle through target
  • 62506 06/08/01 14:15 apar Parity_Start_of_Run_666
  • 62106 06/08/01 13:30 Piotr TARGET NMR
  • 62496 06/08/01 13:29 apar Parity_Start_of_Run_664
  • 62883 06/08/01 13:04 xfzhu problem with laser #4
  • 62488 06/08/01 13:02 LeRose Half wave plate out
  • 62484 06/08/01 12:40 apar Parity_Start_of_Run_662
  • 62477 06/08/01 12:15 apar Parity_Start_of_Run_661
  • 62474 06/08/01 11:50 LeRose Beam is back
  • 62472 06/08/01 11:50 apar Parity_Start_of_Run_660
  • 62470 06/08/01 11:12 LeRose Injector Update
  • 62469 06/08/01 11:09 Zolnierczuk TARGET: NMR
  • 62468 06/08/01 10:35 LeRose Injector Problems
  • 62467 06/08/01 10:02 LeRose Beam Off
  • 62463 06/08/01 09:38 apar Parity_Start_of_Run_659
  • 62455 06/08/01 08:36 Ole Hansen Wrong timestamps in helicity.adl
  • 62847 06/08/01 08:20 R. Michaels chkfeedback script
  • 62057 06/08/01 07:55 kramerk Raster problem solved
  • 62444 06/08/01 07:44 R. Michaels T1 miscounted on L-arm ?
  • 62441 06/08/01 07:39 apar Parity_Start_of_Run_656
  • 62440 06/08/01 07:08 kramerk Target-NMR
  • 62439 06/08/01 07:08 kramerk Target-Field Rotated to 0
  • 62433 06/08/01 06:06 apar Parity_Start_of_Run_655
  • 62431 06/08/01 05:45 R. Michaels another way to run parity analysis
  • 62427 06/08/01 05:25 R. Michaels charge asymmetry feedback results
  • 62426 06/08/01 05:21 kramerk Target-Rotated field to 270
  • 62420 06/08/01 04:20 apar Parity_Start_of_Run_653
  • 62415 06/08/01 03:55 apar Parity_Start_of_Run_652
  • 62412 06/08/01 03:44 R. Michaels RunStart comm. wrong - run summary here
  • 62410 06/08/01 03:34 apar Parity_Start_of_Run_651
  • 62406 06/08/01 03:31 apar Parity_Start_of_Run_650
  • 62401 06/08/01 02:53 kramerk 3He Ref. Cell Pressures
  • 62393 06/08/01 02:37 apar Parity_Start_of_Run_648
  • 62388 06/08/01 02:20 apar Parity_Start_of_Run_647
  • Thursday

  • 62384 06/07/01 19:25 K. McCormick Spectrometer entrance window cover on
  • 62383 06/07/01 19:23 K. McCormick Access to check target lifter
  • 62382 06/07/01 19:12 X. Zheng laser power from May 30 test
  • 62381 06/07/01 19:01 X. Zheng Correction to rate/time of 3He elastic asymmetry measurement
  • 62379 06/07/01 17:00 X. Zheng 3He elastic rate from ref.cell run #1438
  • 62378 06/07/01 16:04 gilman shift summary
  • 62376 06/07/01 16:02 gilman beam!
  • 62375 06/07/01 15:48 Piotr TARGET NMR
  • 62374 06/07/01 15:28 gilman beam! NOT!
  • 62372 06/07/01 14:54 gilman beam coming back
  • 62368 06/07/01 12:59 xfzhu right arm shower/preshower calibration
  • 62367 06/07/01 12:55 gilman beam status
  • 62361 06/07/01 12:12 gilman beam coming back
  • 62360 06/07/01 11:29 apar Parity_Start_of_Run_645
  • 62354 06/07/01 11:21 R. Michaels scaler gate fixed
  • 61965 06/07/01 11:15 averett beam position summary
  • 62356 06/07/01 11:10 gilman controlled access ending
  • 62355 06/07/01 11:07 J.P./Piotr TARGET: He3 reference
  • 62632 06/07/01 10:29 R. Michaels tune status, eP ion chambers
  • 62627 06/07/01 10:28 Kelly file split on 1491
  • 62340 06/07/01 09:44 apar Parity_Start_of_Run_643
  • 62337 06/07/01 09:42 apar Parity_Start_of_Run_642
  • 62333 06/07/01 09:39 apar Parity_Start_of_Run_641
  • 62328 06/07/01 09:30 gilman controlled access to look at right arm dead time
  • 62327 06/07/01 09:09 R. Michaels re: 150% live-time
  • 62322 06/07/01 08:39 apar Parity_Start_of_Run_639
  • 62713 06/07/01 08:00 J. Roche shift summary
  • 62319 06/07/01 08:00 Kelly shift summary
  • 61926 06/07/01 07:15 Nilanga OTR with raster on
  • 62317 06/07/01 06:43 kramerk 3He Ref. Cell Problem Hypothesis
  • 62313 06/07/01 06:25 apar Parity_Start_of_Run_638
  • 62307 06/07/01 06:03 apar Parity_Start_of_Run_637
  • 62305 06/07/01 05:47 kramerk Problems with 3He gas fill
  • 62302 06/07/01 05:36 apar Parity_Start_of_Run_636
  • 62301 06/07/01 05:27 Armstrong raster check after Moller
  • 62298 06/07/01 05:19 apar Parity_Start_of_Run_635
  • 62295 06/07/01 05:09 apar Parity_Start_of_Run_634
  • 62294 06/07/01 04:38 Eugene Moller results (preliminary)
  • 62293 06/07/01 04:01 X. Zheng false asymmetry study - right arm
  • 62292 06/07/01 04:01 X. Zheng false asymmetry study - left arm
  • 62287 06/07/01 00:22 kramerk Ref. Cell. Evacuated
  • 62282 06/07/01 00:06 jpchen Swing Shift summary
  • Wednesday

  • 61996 06/06/01 23:57 X. Zheng Gas cherenkov detector calibration
  • 62279 06/06/01 23:56 apar Parity_Start_of_Run_633
  • 62275 06/06/01 23:43 apar Parity_Start_of_Run_632
  • 62271 06/06/01 23:36 X. Zheng June 06 target polarization from EPR
  • 62266 06/06/01 23:12 Higinbotham Mini-Current Calibration
  • 62265 06/06/01 23:08 atrig Trigger-Download
  • 62263 06/06/01 23:05 jpchen/X. Zh EPR measurement
  • 62259 06/06/01 22:51 apar Parity_Start_of_Run_630
  • 62258 06/06/01 22:49 atrig Trigger-Download
  • 62253 06/06/01 22:37 Antonin Vach Parity DAQ analyzer is working !
  • 62250 06/06/01 22:29 atrig Trigger-Download
  • 62249 06/06/01 22:27 atrig Trigger-Download
  • 62248 06/06/01 22:27 Antonin Charge Asymmetry runs for feedbacks
  • 61958 06/06/01 22:11 Nilanga Bull's eye (0, +2)
  • 62241 06/06/01 22:09 jpchen completed charge asymmetry slope and 12C false asymmetry runs
  • 62236 06/06/01 21:57 apar Parity_Start_of_Run_629
  • 62229 06/06/01 21:47 apar Parity_Start_of_Run_628
  • 62223 06/06/01 21:41 apar Parity_Start_of_Run_627
  • 62216 06/06/01 21:29 apar Parity_Start_of_Run_625
  • 61819 06/06/01 21:21 R. Michaels more on raster
  • 62209 06/06/01 21:20 apar Parity_Start_of_Run_624
  • 62202 06/06/01 21:07 apar Parity_Start_of_Run_622
  • 61800 06/06/01 20:54 X.Zheng Right arm gas cherenkov raw ADCs
  • 62582 06/06/01 20:49 apar Parity_Start_of_Run_683
  • 62579 06/06/01 20:49 apar Parity_Start_of_Run_682
  • 62183 06/06/01 20:35 jpchen Beam back (RF problem not fixed yet).
  • 62180 06/06/01 20:29 apar Parity_Start_of_Run_618
  • 62176 06/06/01 20:27 apar Parity_Start_of_Run_617
  • 62174 06/06/01 20:10 reitz right trigger
  • 62173 06/06/01 20:01 jpchen right arm deadtime problem
  • 62172 06/06/01 19:43 jpchen Beam off
  • 62161 06/06/01 19:05 apar Parity_Start_of_Run_615
  • 61769 06/06/01 18:57 A. Deur back to no collimators
  • 62157 06/06/01 18:53 apar Parity_Start_of_Run_614
  • 62155 06/06/01 18:52 apar Parity_Start_of_Run_613
  • 62153 06/06/01 18:49 jpchen charge asymmetry curve measurement
  • 62439 06/06/01 18:49 kramerk Target-Field Rotated to 0
  • 62151 06/06/01 18:48 apar Parity_Start_of_Run_612
  • 62150 06/06/01 18:42 jpchen raster set to 4mm and beam to 9 uA on 12C
  • 62145 06/06/01 18:23 apar Parity_Start_of_Run_611
  • 62140 06/06/01 18:07 jpchen Beam back
  • 62139 06/06/01 18:07 apar Parity_Start_of_Run_610
  • 62138 06/06/01 17:24 X. Zheng June 05 Target polarization from EPR
  • 62135 06/06/01 16:50 Antonin Vach Pedestal runs for Parity ADCs
  • 62134 06/06/01 16:34 Antonin/Piot Parity DAQ: halog entry script
  • 62133 06/06/01 16:27 apar Parity_Start_of_Run_608
  • 62132 06/06/01 16:24 apar TEST
  • 61857 06/06/01 16:20 R. Michaels problem of R-arm scalers (fixed)
  • 61855 06/06/01 16:17 A. Deur current distribution for the raster
  • 62119 06/06/01 15:38 Ole Hansen Reposition beam at BPMA to minimize beam trajectory angle through target
  • 62108 06/06/01 14:31 Ole Hansen Centered beam on BeO
  • 62107 06/06/01 14:18 Piotr TARGET: NMR
  • 62106 06/06/01 12:08 Piotr TARGET NMR
  • 62496 06/06/01 11:47 apar Parity_Start_of_Run_664
  • 62384 06/06/01 09:29 K. McCormick Spectrometer entrance window cover on
  • 62098 06/06/01 09:04 Piotr Target: filled with N2
  • 62097 06/06/01 08:20 kramerk Shift Summary
  • 62093 06/06/01 07:43 kramerk ROC15 rebooted and run numbers out of sync
  • 62067 06/06/01 04:58 kramerk Reference Cell N2 5 atm
  • 62056 06/06/01 04:03 Kelly run 1389
  • 62050 06/06/01 03:34 Z.-E. Mezian empty target run
  • 62048 06/06/01 03:16 Higinbotham ARC Energy (BDL Only)
  • 62045 06/06/01 02:51 Kelly raster and beam position
  • 61765 06/06/01 01:08 A. Deur coda
  • 62038 06/06/01 00:43 Piotr DAQ: runend.tcl (fixes)
  • Tuesday

  • 62035 06/05/01 23:51 Lingyan swing shift summary
  • 62305 06/05/01 22:18 kramerk Problems with 3He gas fill
  • 62001 06/05/01 19:26 jpchen Reference cell inspection
  • 62000 06/05/01 19:19 R. Michaels/ raster setup
  • 62282 06/05/01 18:56 jpchen Swing Shift summary
  • 61996 06/05/01 18:54 X. Zheng Gas cherenkov detector calibration
  • 62279 06/05/01 18:50 apar Parity_Start_of_Run_633
  • 61985 06/05/01 15:36 R. Michaels N2 ref cell spot
  • 62372 06/05/01 15:23 gilman beam coming back
  • 61982 06/05/01 15:12 Nilanga something strange with reference cell
  • 62265 06/05/01 14:22 atrig Trigger-Download
  • 61979 06/05/01 14:20 averett SPOT for ref. cell runs
  • 61971 06/05/01 13:44 Nilanga harp with raster on
  • 61965 06/05/01 13:18 averett beam position summary
  • 61963 06/05/01 12:49 Nilanga Bull's eye (0, -2)
  • 61962 06/05/01 12:48 averett bullseye 0,-2
  • 61958 06/05/01 12:36 Nilanga Bull's eye (0, +2)
  • 61685 06/05/01 12:28 A. Deur coda and run 1319
  • 61682 06/05/01 12:26 D. Lhuillier Compton Polarimeter
  • 61950 06/05/01 12:13 averett bullseye -2,0
  • 61947 06/05/01 12:08 Nilanga Bull's eye (2,0)
  • 61939 06/05/01 11:49 Nilanga Bull';s eye scan (0,0)
  • 61938 06/05/01 11:48 averett bullseye scan
  • 61937 06/05/01 11:44 averett DAQ synch
  • 61929 06/05/01 11:17 averett DAQ synch unplugged
  • 61927 06/05/01 10:44 Nilanga dose rates
  • 61925 06/05/01 10:40 NIlanga Harp scan with raster on
  • 61922 06/05/01 10:09 Higinbotham Check of BPM numbers
  • 61921 06/05/01 10:05 Nilanga OTR
  • 61902 06/05/01 09:21 E.Chudakov Harp scan
  • 61901 06/05/01 08:38 R. Michaels Charge Asymmetry Tests
  • 62172 06/05/01 08:12 jpchen Beam off
  • 62153 06/05/01 06:59 jpchen charge asymmetry curve measurement
  • 61861 06/05/01 05:29 E.Chudakov Harp scan
  • 61860 06/05/01 04:11 E.Chudakov Pulsed beam appears
  • 61859 06/05/01 00:47 P.Zolnierczu DAQ: runend.tcl charge asymmetry problem (fixed)
  • Monday

  • 61587 06/04/01 23:26 busato signal on channel 6. cosmic run
  • 61585 06/04/01 23:02 R. Michaels fastbus helicity flag checks
  • 61855 06/04/01 22:58 A. Deur current distribution for the raster
  • 61854 06/04/01 22:55 Roedelbronn current calibration
  • 61843 06/04/01 21:50 Roedelbronn Beam returning...
  • 61842 06/04/01 20:54 Roedelbronn Beam delay
  • 61837 06/04/01 20:00 Roedelbronn beam
  • 62107 06/04/01 17:46 Piotr TARGET: NMR
  • 61835 06/04/01 17:40 xfzhu BeO Camera
  • 0 06/04/01 17:28 no_name no_key
  • 61832 06/04/01 16:39 Higinbotham ARC Scanner Fixed
  • 61831 06/04/01 15:41 jpchen Beam by 8 pm.
  • 61824 06/04/01 15:06 Higinbotham MEDM BPM Offset Screen
  • 61823 06/04/01 14:38 Cates/Zolnie TARGET: NMR/EPR
  • 62216 06/04/01 13:54 apar Parity_Start_of_Run_625
  • 61820 06/04/01 12:38 jpchen BeO camara
  • 61819 06/04/01 12:25 R. Michaels more on raster
  • 61818 06/04/01 12:14 korsch Life Time in HRS-R
  • 61815 06/04/01 11:24 R. Michaels re: raster and coda
  • 61814 06/04/01 11:06 X. Jiang checked helicity-sign raw adc in espace
  • 61811 06/04/01 10:52 X. Zheng ADC software cut in espace
  • 61808 06/04/01 10:38 X. Jiang left-arm gas adc, no limit on ch4096.
  • 61806 06/04/01 10:18 X. Zheng Right Arm Shower/Preshower performance
  • 61803 06/04/01 10:11 X. Zheng General performance of gas cherenkov detectors
  • 61800 06/04/01 08:28 X.Zheng Right arm gas cherenkov raw ADCs
  • 61797 06/04/01 08:08 A. Deur more on the raster
  • 61796 06/04/01 07:31 A. Deur coda
  • 61795 06/04/01 07:27 A. Deur raster problem.
  • 61793 06/04/01 07:09 kramerk Target - Results from NMR/EPR Tests
  • 61790 06/04/01 07:05 A. Deur end of optics runs. begin N2 ref cell runs.
  • 61784 06/04/01 06:49 X.Zheng Target polarization from EPR
  • 61776 06/04/01 06:39 A. Deur radiation level
  • 61765 06/04/01 06:27 A. Deur coda
  • 61762 06/04/01 06:22 A. Deur raster run
  • 61758 06/04/01 05:46 A, Deur Momentum change
  • 61753 06/04/01 05:33 A. Deur Sieve slit in
  • 61740 06/04/01 05:04 A. Deur polarimetry done. No OTR. Start optic
  • 61738 06/04/01 04:07 A. Deur target polarimetry
  • 61737 06/04/01 04:06 A. Deur beam position scan
  • 61985 06/04/01 03:21 R. Michaels N2 ref cell spot
  • 62372 06/04/01 03:19 gilman beam coming back
  • 61710 06/04/01 03:04 A. Deur coda problem
  • 61708 06/04/01 02:59 busato about ADCSUM
  • 61706 06/04/01 02:42 busato extensive analysis of the run 20317
  • 61695 06/04/01 02:03 A. Deur Compton runs and empty target
  • 61694 06/04/01 01:51 busato extensive analysis of the run 1316
  • 61693 06/04/01 01:48 V. Sulkosky Junk Runs with EP Target
  • 61690 06/04/01 01:38 A. Deur Beam back. Start new eP run
  • 61687 06/04/01 01:25 A. Deur beam down 20 min
  • 61686 06/04/01 01:23 A. Deur How long the runs should be ?
  • 61685 06/04/01 01:13 A. Deur coda and run 1319
  • 61682 06/04/01 01:06 D. Lhuillier Compton Polarimeter
  • 61950 06/04/01 00:51 averett bullseye -2,0
  • 61673 06/04/01 00:18 Lingyan bpmb offset
  • 61672 06/04/01 00:07 Lingyan swing shift summary
  • Sunday

  • 61668 06/03/01 23:39 A. Deur number of track for VDC
  • 61667 06/03/01 23:35 Sarah Harp after Beam Alignment
  • 61657 06/03/01 22:42 xfzhu BeO camera is ok now
  • 61656 06/03/01 22:20 Higinbotham BeO / Carbon Foil Target
  • 61655 06/03/01 21:51 Lingyan access to move the camera
  • 61652 06/03/01 20:33 Lingyan harp
  • 61649 06/03/01 19:52 jpchen HARP problem fixed
  • 61643 06/03/01 15:59 R. Michaels problem of S1-L-1 TDC fixed
  • 61635 06/03/01 13:40 X. Jiang checked helicity in espace.
  • 61627 06/03/01 07:34 X. Zheng target ioc reboot
  • 61625 06/03/01 04:56 kramerk FSD Problem
  • 61624 06/03/01 04:38 kramerk Target - Rebooted IOC
  • 61623 06/03/01 04:31 J. Roche magnets set up
  • 61622 06/03/01 04:24 kramerk Target - Target IOC Reboot
  • 62000 06/03/01 02:36 R. Michaels/ raster setup
  • Saturday

  • 61614 06/02/01 20:59 K. McCormick EPICS control instructions for He3 target
  • Friday

  • 61587 06/01/01 12:23 busato signal on channel 6. cosmic run
  • 61585 06/01/01 12:09 R. Michaels fastbus helicity flag checks